OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

oops yes.

OpenStudy (anonymous):

Can you give the options?

OpenStudy (anonymous):

use substitution or elimination

OpenStudy (anonymous):

−2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

ok

OpenStudy (luigi0210):

Solve the first one x and then substitute that's make things so much easier. And hey @Peasant1 !

OpenStudy (campbell_st):

so you need to eliminate y multiple the 1st equation by 5 and the 2nd equation by -6 and you get 15x - 30y = 15 - 42x + 30y =66 add the equations to eliminate y -27x = 81 I'll let you solve for x

OpenStudy (anonymous):

Howdy! @Luigi0210

OpenStudy (anonymous):

@campbell_st Is right.

OpenStudy (anonymous):

So is @Luigi0210

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No problem.

OpenStudy (luigi0210):

What was your final answer?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

x=-3....

OpenStudy (anonymous):

-42x+30y=66 15x-30y=15...........by solving......-27x=81........=.....x=-3

OpenStudy (anonymous):

it was already solved lol but thanks anyway